OpenStudy (anonymous):

3. When solving equations with fractions we apply the same operations as we do when we solve with integers, so what is the first step used to solve the following equations? b - 3/5 = 1/5 A. subtract 3/5 B. take inverse of 3/5 C. move x to the other side D. add 3/5

OpenStudy (anonymous):

D. add 3/5 to both sides

OpenStudy (anonymous):

4. On Solving Equations with Fractions: Adding or Subtracting Part 2 page, what property is used to solve example 1? x + 1/3 = 2/3 A. Addition Property of Equality B. Inverse Property of Equality C. Associative Property of Equality D. Subtraction Property of Equality

OpenStudy (anonymous):

D. Subtraction Property of Equality

OpenStudy (anonymous):

subtract 1/3 from both sides

OpenStudy (anonymous):

thank you .

OpenStudy (anonymous):

1. What is the inverse operation used to solve the following equation found on Solving Equations with Fractions page? t-3 2/3=1 1/3 A. add 3 2/3 B. Subtract t C. subtract 3 2/3 D. add 1 1/3

OpenStudy (anonymous):

Search on the net and read a fes solved examples and u will understand what is happening...

OpenStudy (anonymous):

A. add 3 2/3

OpenStudy (anonymous):

thankyou i can now plug them in in the problems and see how theyre answered